OpenStudy (anonymous):

please help i will reward 2 medals

OpenStudy (anonymous):

What do you need help with?

OpenStudy (anonymous):

ok A line passes through (3, 7) and (6, 9). Which equation best represents the line? y = 3 over 2x + 5 y = 2 over 3x + 5 y = 3x + 2 y = 2 over 3x + 2

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

Did you try substituting the points in the equations?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

but i dont get how to do that

OpenStudy (anonymous):

Alright, I'll explain.

OpenStudy (anonymous):

thanks btw

OpenStudy (anonymous):

so @ShizukaTheOtaku

OpenStudy (theraggedydoctor):

You can't give out two medals, btw..

OpenStudy (anonymous):

yes i can

OpenStudy (theraggedydoctor):

No you can't.

OpenStudy (theraggedydoctor):

You can only give out one.

OpenStudy (anonymous):

i can start a new convo with nothing

OpenStudy (anonymous):

I'll do the first for you. So 3 is the x value, and 7 is the y. y = 3/ (2x + 5) which is (7) = 3/ (2(3) +5) So as you can see we are basically substituting the numbers in the place of x and y. And to solve... 7 = 3 /11 We do the same for the other coordinate (6,9) 9 = 3 (2(6) + 5) 9 = 51 They aren't the same. Both sides of the equation must be equal.

OpenStudy (anonymous):

Those two points aren't on the same line.

OpenStudy (anonymous):

ok so its b? if i am understanding @ShizukaTheOtaku

OpenStudy (anonymous):

Right @ShizukaTheOtaku ??

OpenStudy (anonymous):

How did you get that answer?

OpenStudy (anonymous):

i did what you said and i got b

OpenStudy (anonymous):

i am guessing it is wrong

OpenStudy (anonymous):

By over do you mean 3 / (2x + 5) ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Actually I'm completely stumped now because the rule I know doesn't seem to apply here. Maybe you can tag someone else who knows? Sorry for not being able to help.

OpenStudy (anonymous):

np u tried

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

@samanthagreer

OpenStudy (anonymous):

A line passes through (3, 7) and (6, 9). Which equation best represents the line? y = 3 over 2x + 5 y = 2 over 3x + 5 y = 3x + 2 y = 2 over 3x + 2

OpenStudy (anonymous):

over means /

OpenStudy (anonymous):

so no one knows what to do?

OpenStudy (anonymous):

so no one knows what to do?

OpenStudy (welshfella):

the way you have written the equations is a little confusing- for example the first one which of the following is correct:-